# MVGT: A Multi-view Graph Transformer Based on Spatial Relations for EEG Emotion Recognition

### v1

@@ IV-D Implementation Details of MVGT
$\bm{X^{\prime}}$ $=\text{GraphNorm}(\bm{X}),$ (9)
$\bm{H}^{(0)}$ $=\text{SpatialEncoding}+\text{Proj}(\bm{X^{\prime}}),$ (10)
@@
$\bm{H^{\prime}}^{(l)}$ $=\text{MHA}(\text{LN}(\bm{H}^{(l-1)}))+\bm{H}^{(l-1)},$ (11)
$\bm{H}^{(l)}$ $=\text{FFN}(\text{LN}(\bm{H^{\prime}}^{(l)}))+\bm{H^{\prime}}^{(l)},$ (12)
@@ V-A Datasets
For our experiments, we selected the SEED[5] and SEED-IV[47] datasets to evaluate the effectiveness of our model. These datasets consist of EEG signals recorded from subjects while they watched emotion-eliciting videos.
SEED dataset comprises data from 15 subjects who participated in three sessions, each separated by at least one week. Each sessions consists of 15 trials capturing emotional labels, with the emotion labels being positive, negative, and neutral.
SEED-IV dataset is constituted by EEG signals from 15 subjects across three separate sessions conducted at different times, using the same device as the SEED dataset. This dataset encompasses four emotion labels: neutral, sad, fear, and happy. In each session, each subject underwent 24 trials.
@@ V-B Settings
V-C Baseline Models
  • $\bullet$
    DGCNN[12]: A dynamic graph neural network method based on Chebyshev polynomials dynamically learns inter-channel relations in emotion recognition.
  • $\bullet$
    BiHDM[10]: This model employs a pairwise subnetwork to capture the discrepancy between the left and right hemispheres of the brain.
  • $\bullet$
    R2G-STNN[11]: A model that captures spatial-temporal features from local to global scales for emotion classification.
  • $\bullet$
    RGNN[13]: A regularized GNN that learns topological relationships between channels.
  • $\bullet$
    MD-AGCN[14]: An adaptive GNN that comprehensively considers temporal domain, frequency domain, and brain functional connectivity.
  • $\bullet$
    MV-SSTMA[31]: A multi-view masked autoencoder combining CNN and Transformer for emotion recognition.
  • $\bullet$
    EmoGT[16]: A elastic graph Transformer network that integrates temporal and spatial information.
@@ V-C Baseline Models
V-D Results Analysis
We compare the classification results based on the SEED and SEED-IV datasets with recent state-of-the-art models, as shown in Table 1. It is evident that our proposed model significantly outperforms the baseline models under the same experimental settings. In the experiments on the SEED dataset, the model adopting the FRONTAL scheme achieved the best performance, with a classification accuracy of 96.45%. The LOBE scheme also achieved a slightly superior accuracy of 95.36%, compared to other models. For the SEED-IV dataset, the classification accuracy under the GENERAL scheme was 93.57%, reaching the best performance compared to baseline models. The MVGT model also achieved commendable results under other division schemes. Overall, our model achieved the best recognition accuracy compared to the baselines. The results suggest that selecting the specific division scheme relevant to the emotion task could enhance the expressive power of MVGT.
(a) SEED
(b) SEED-IV
@@ V-D Results Analysis
Fig. 33(a) and 33(b) illustrate the confusion matrices of MVGT-F on the SEED and MVGT-G on the SEED-IV, respectively. The values represent the recognition accuracy of the model for different emotion classes. For the SEED dataset, our model achieved the highest accuracy in recognizing positive emotions (98.12%), followed by neutral emotions (96.38%), with negative emotions being slightly lower (94.73%). Only 0.33% of positive emotion samples were misclassified as negative, while only 0.76% of negative emotion samples were recognized as positive, indicating the model’s effectiveness in distinguishing valence changes. For the SEED-IV dataset, our model performed best in recognizing neutral emotions, with an accuracy of 95.90%, while its performance on happy emotions was slightly lower than the other three emotions, with an accuracy of 90.76%. This could be attributed to the GENERAL scheme setting, making the model more sensitive to balanced emotions.
V-E Ablation Study
To validate the effectiveness of spatial encodings, we conducted ablation experiments on the SEED and SEED-IV datasets, as presented in Table II. By removing both types of spatial encoding, we repeated the aforementioned experiments under the same experimental settings. On the SEED dataset, the model achieved an accuracy of 93.79% with a standard deviation of 7.15%. Compared to MVGT-F, the accuracy decreased by 2.66% and the standard deviation increased by 2.75% after removing spatial encodings. For the SEED-IV dataset, the accuracy dropped by 4.08%, resulting in 89.49%, with the standard deviation rising by 1.80% to 10.40%, when compared to MVGT-G. The experiments demonstrate that incorporating spatial structure information benefits the model performance in emotion recognition tasks. Under experimental settings that consider only geometric structure or brain region structure, the model’s classification accuracy improved over the plain model without any spatial encoding.
SEED
[Uncaptioned image][Uncaptioned image]
[Uncaptioned image][Uncaptioned image]
[Uncaptioned image][Uncaptioned image]
Refer to captionRefer to caption
SEED-IV
Refer to captionRefer to caption
Layer-1
Refer to captionRefer to caption
Layer-2
Refer to captionRefer to caption
Layer-3
Refer to captionRefer to caption
Layer-4
@@ V-E Ablation Study
Evidently, when considering both types of spatial structures simultaneously, the model performance significantly surpassed that of the plain model and models using only single spatial information. This indicates the effectiveness of our proposed spatial encodings and confirms that the expressive power of the Graph Transformer relies on the spatial structure and positional encoding.
V-F Visualization of Inter-channel relations
@@ VI Conclusions
In future work, we will focus on the following aspects: (1) designing more optimal structural encodings, such as data-driven methods for adaptive structural encoding; (2) attempting to combine various handcrafted features and exploring the possibility of extracting effective EEG features through neural networks; (3) investigating emotion recognition methods based on multimodal physiological signals.

### v2

@@ IV-D Implementation Details of MVGT
$\bm{X^{\prime}}_{s}$ $=\text{GraphNorm}(\bm{X}_{s}),$ (10)
$\bm{H}^{0}_{s}$ $=\text{SpatialEncoding}+\text{Proj}(\bm{X^{\prime}}_{s}),$ (11)
@@
$\bm{H^{\prime}}^{l}_{s}$ $=\text{MHA}(\text{LN}(\bm{H}^{l-1}_{s}))+\bm{H}^{l-1}_{s},$ (12)
$\bm{H}^{l}_{s}$ $=\text{FFN}(\text{LN}(\bm{H^{\prime}}^{l}_{s}))+\bm{H^{\prime}}^{l}_{s},$ (13)
@@ V-A Datasets
The datasets consist of EEG signals recorded from subjects while they watched emotion-eliciting videos.
SEED [5] dataset comprises data from 15 subjects who participated in three sessions, each separated by at least one week. Each sessions consists of 15 trials capturing emotional labels, with the emotion labels being positive, negative, and neutral.
SEED-IV [47] dataset is constituted by EEG signals from 15 subjects across three separate sessions conducted at different times, using the same device as the SEED dataset. This dataset encompasses four emotion labels: neutral, sad, fear, and happy. In each session, each subject underwent 24 trials.
@@ V-B Settings
V-C Results Analysis
We compare the classification results based on the SEED and SEED-IV datasets with recent state-of-the-art models,
@@ V-C Results Analysis
as shown in Table 1. It is evident that our proposed model significantly outperforms the baseline models under the same experimental settings. In the experiments on the SEED dataset, the model adopting the FRONTAL scheme achieved the best performance, with a classification accuracy of 96.45%. The LOBE scheme also achieved a slightly superior accuracy of 95.36%, compared to other models. For the SEED-IV dataset, the classification accuracy under the GENERAL scheme was 93.57%, reaching the best performance compared to baseline models. The MVGT model also achieved commendable results under other division schemes. Overall, our model achieved the best recognition accuracy compared to the baselines. The results suggest that selecting the specific division scheme relevant to the emotion task could enhance the expressive power of MVGT.
@@
Fig. 3 illustrates the confusion matrices of MVGT-F on the SEED and MVGT-G on the SEED-IV, respectively. The values represent the recognition accuracy of the model for different emotion classes. For the SEED dataset, our model achieved the highest accuracy in recognizing positive emotions (98.12%), followed by neutral emotions (96.38%), with negative emotions being slightly lower (94.73%). Only 0.33% of positive emotion samples were misclassified as negative, while only 0.76% of negative emotion samples were recognized as positive, indicating the model’s effectiveness in distinguishing valence changes. For the SEED-IV dataset, our model performed best in recognizing neutral emotions, with an accuracy of 95.90%, while its performance on happy emotions was slightly lower than the other three emotions, with an accuracy of 90.76%. This could be attributed to the GENERAL scheme setting, making the model more sensitive to balanced emotions.
@@
V-D Ablation Study
To validate the effectiveness of spatial encodings, we conducted ablation experiments on the SEED and SEED-IV datasets, as presented in Table II. By removing both types of spatial encoding, we repeated the aforementioned experiments under the same experimental settings. On the SEED dataset, the model achieved an accuracy of 93.79% with a standard deviation of 7.15%. Compared to MVGT-F, the accuracy decreased by 2.66% and the standard deviation increased by 2.75% after removing spatial encodings. For the SEED-IV dataset, the accuracy dropped by 4.08%, resulting in 89.49%, with the standard deviation rising by 1.80% to 10.40%, when compared to MVGT-G. The experiments demonstrate that incorporating spatial structure information benefits the model performance in emotion recognition tasks. Under experimental settings that consider only geometric structure or brain region structure, the model’s classification accuracy improved over the plain model without any spatial encoding. Evidently, when considering both types of spatial structures simultaneously, the model performance significantly surpassed that of the plain model and models using only single spatial information. This indicates the effectiveness of our proposed spatial encodings and confirms that the expressive power of the Graph Transformer relies on the spatial structure and positional encoding.
@@ V-D Ablation Study
V-E Visualization of Inter-channel relations